OpenStudy (anonymous):

x + 7 = -2x(4x + 10) Not sure exactly how to do this, start it..

OpenStudy (anonymous):

x=-7/29

OpenStudy (anonymous):

wait hold on minor error

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Hi, I would start by multiplying everuthing inside the brackets by -2x, giving x + 7 = -8x^2-20x

OpenStudy (anonymous):

Would i have to 8x^2 by 20 x? next

OpenStudy (anonymous):

once you do that you put everything equal to zero. it should be 8x^2 +21x+7=0

OpenStudy (anonymous):

then you can factor. to find x